OpenStudy (anonymous):

If f(c) = 0, which of the following statements must be true? (Points : 2) x + c is a factor of f(x). The point (0,c) lies on the graph of f(x). x - c divides evenly into f(x). The point (-c,0) does not lie on the graph of f(x).

OpenStudy (nincompoop):

understand that f(c) = 0 is another way of indicating y=0

OpenStudy (anonymous):

hmm .... it cannot be x+c because the format is x-c the point 0,c makes sense x-c divides evenly also makes sense so would it be one of the middle two?

OpenStudy (nincompoop):

is y dependent or independent?

OpenStudy (anonymous):

independent

OpenStudy (nincompoop):

y is always dependent… it is dependent on the values of x or functions associated with x

OpenStudy (anonymous):

opps... okay

OpenStudy (nincompoop):

let me try to give you some examples of what I mean let us say I have a f(x) = 2x^2 as I have said before, it is another way of indicating y = 2x^2 so when we evaluate f(-2) it will be y = 2(-2)^2 now the value of y rest solely on the values of x or functions associated with x it will be y = 2*4 y = 8 or f(-2) = 8

OpenStudy (anonymous):

then it would be -c,0 does not lie on the graph?

OpenStudy (nincompoop):

let me try another function say: g(x) = 1/x and we were to evaluate g(4) and then g(f(x)) first we do this, g(4) = 1/x g(4) = 1/4 or simply y = 1/4 now let us evaluate g(f(x)) first we obtain f(x) from our previous example f(x) = 2x^2 then evaluate g(f(x)) it will be g(f(x)) 1/(2x^2) since our f(x) is now the x in the g(x)

OpenStudy (nincompoop):

in your question, we are clearly missing some information based on the options of answers provided. perhaps your question is relevant to another set of question(s) or previous question. if not, then every time we see f(x) = 0 it must clearly state that y = 0 the format maybe tricky for starters, but it is simply telling you f(x) = 0*x and no matter what we do, it will always come back as zero.

OpenStudy (nincompoop):

are you following me?

OpenStudy (anonymous):

not really... i got lost with your last post

OpenStudy (nincompoop):

did you follow before the last post?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

okay, good

OpenStudy (nincompoop):

what does f(c) = 0 mean then?

OpenStudy (anonymous):

um if im correct i think its x-c divides evenly into f(c) right?

OpenStudy (anonymous):

*f(x)

OpenStudy (nincompoop):

where are you getting the x?

OpenStudy (anonymous):

the independent variable?

OpenStudy (nincompoop):

but c is our independent variable in f(c) = 0 you can replace x with anything in a function it is just a place holder for any independent variable f(monkey) = 0 is just another way of saying f(monkey) = zero times monkey

OpenStudy (anonymous):

mkay

OpenStudy (nincompoop):

that is why I said earlier that you might be missing some information or have forgotten to include something else that makes x okay to exist

OpenStudy (anonymous):

ohh okay!! :) thank you!

OpenStudy (nincompoop):

there is no x in f(c) = 0